OpenStudy (anonymous):

MEDAL + FAN 2 to the 3rd plus 3 to the 3rd power

Mehek (mehek14):

\(\Large{2^3 + 3^3?}\)

Mehek (mehek14):

like this?

OpenStudy (anonymous):

YESS @Mehek14

OpenStudy (anonymous):

2^3= 8 3^3=27 so, 8+27= 35